### v1

# Q-Net: Transferable Queue Length Estimation via Kalman-based Neural Networks

Abstract
Estimating queue lengths at signalized intersections remains a challenge in traffic management, especially under partially observed conditions where vehicle flows are not fully captured. This paper introduces Q-Net, a data-efficient and interpretable framework for queue length estimation that performs robustly even when traffic conservation assumptions are violated. Q-Net integrates two widely available and privacy-friendly data sources: (i) vehicle counts from loop detectors near stop lines, and (ii) aggregated floating car data (aFCD), which divides each road section into segments and provides segment-wise average speed measurements. These data sources often differ in spatial and temporal resolution, creating fusion challenges. Q-Net addresses this by employing a tailored state-space model and an AI-augmented Kalman filter, KalmanNet, which learns the Kalman gain from data without requiring prior knowledge of noise covariances or full system dynamics. We build on the vanilla KalmanNet pipeline to decouple measurement dimensionality from section length, enabling spatial transferability across road segments. Unlike black-box models, Q-Net maintains physical interpretability, with internal variables linked to real-world traffic dynamics. Evaluations on main roads in Rotterdam, the Netherlands, demonstrate that Q-Net outperforms baseline methods by over 60% in Root Mean Square Error (RMSE), accurately tracking queue formation and dissipation while correcting aFCD-induced delays. Q-Net also demonstrates strong spatial and temporal transferability, enabling deployment without costly sensing infrastructure like cameras or radar. Additionally, we propose a real-time variant of Q-Net, highlighting its potential for integration into dynamic, queue-based traffic control systems.
1 Introduction
Traffic queues raise operational, safety, and environmental concerns that necessitate effective monitoring and control systems. Congestion in cities reduces network throughput and limits agglomeration economies (Sweet, 2011). Beyond efficiency concerns, long queues increase accident likelihood (Pasquale et al., 2018), elevate noise levels (Chevallier et al., 2009), contribute to driver stress and mental health issues (Nadrian et al., 2019), and generate localized air pollution due to stationary vehicles (Matzoros and Van Vliet, 1992). Real-time queue length estimation addresses these challenges by enabling adaptive traffic control strategies that dynamically optimize signal timing in response to changing conditions, while accurate queue length estimation more broadly supports the development of advanced queue-based control algorithms and network-wide optimization strategies (Li et al., 2021; Hoogendoorn et al., 2016). However, despite being a longstanding focus in traffic management, queue length estimation remains challenging due to incomplete sensor coverage, unpredictable driver behavior, and intricate network topologies.
Queue length estimation methods at signalized intersections are characterized by their type of inputs. These input data can be Lagrangian and Eulerian types. Lagrangian data are collected from mobile sensors that move with traffic flow, primarily through GPS-equipped probe vehicles. Recent research has extensively explored probe vehicle trajectory data for queue length estimation, employing methodologies ranging from shockwave theory to probabilistic likelihood-based approaches (Cheng et al., 2011; Hu et al., 2022; Zhao et al., 2019; Cheng et al., 2012; Zhao et al., 2021). However, these methods typically require prior knowledge of probe vehicle penetration rates and raise privacy concerns, as participants must reveal their positions to external traffic monitoring systems (Rose, 2006).
In contrast, Eulerian data capture traffic conditions over fixed spatial segments, coming from either stationary sensors or aggregated mobile sources. Video-based systems, enhanced by computer vision, can detect queue ends (Umair et al., 2021; Albiol et al., 2011), though their performance may degrade under poor lighting or bad weather. Radar systems, while capable of providing direct queue measurements, are expensive to install and maintain (Bernas et al., 2018). Loop detectors are among the most widely used sensors due to their ability to provide traffic counts, occupancy, and speed measurements at a low cost (MÜCK, 2002; Liu et al., 2009; Geroliminis and Skabardonis, 2011; Abewickrema et al., 2023). Recently, aggregated floating car data (aFCD) from commercial vendors like INRIX, HERE, Be-Mobile, and TomTom represents another Eulerian source. These data are anonymized and aggregated over road segments and time intervals to comply with privacy regulations, offering advantages in terms of data availability without requiring high penetration rates (van der Loop et al., 2017). The aFCD is of widespread use in France (Gitahi et al., 2020), Turkey (Altıntaşı, 2018), and the Netherlands (Mobiliteitsscan, ).
Despite advancements in high-resolution video analytics and individual-level probe vehicle tracking, many cities still rely on conventional data sources, such as loop detectors, due to infrastructure constraints, limited budgets, and privacy regulations. Given these practical considerations, this study focuses on a queue length estimation that uses two widely available data sources as input: loop detector measurements and aFCD, which ensures broad applicability across diverse urban settings and aligns well with infrastructure where these data are readily available.
1.1 Literature review
Queue length estimation at signalized intersections can generally be categorized into four approaches: the input-output models (Webster, 1958; Sharma et al., 2007), shockwave theory (Lighthill and Whitham, 1955; Liu et al., 2009), speed-based methods (Dinh et al., 2014), and filtering algorithms (Yin et al., 2018).
Input-output models estimate queue length based on the conservation equation of cumulative traffic inflow and outflow at signalized sections. Traditionally, researchers have relied on vehicle counts recorded by road detectors to characterize the queuing process (Webster, 1958). Later, Vigos et al. (2008) enhanced this approach by applying real-time flow and occupancy measurements to adjust queue length estimates in accordance with conservation laws. While the input-output model is simple and straightforward, it does not capture the spatiotemporal dynamics of queues and is highly sensitive to the data measurement quality. More recent studies have improved input-output models by integrating probe vehicle data, enabling real-time estimation of the queue tail position (Wang et al., 2021; Tang et al., 2023), but these models require a minimal penetration rate.
Shockwave theory describes the queue formation and dissipation dynamics at signalized intersections, with shockwave speed indicating traffic state transitions. Liu et al. (2009) used loop detector data for cycle-wise queue estimation, later refined by Abewickrema et al. (2023) to handle both oversaturated and undersaturated conditions. These methods rely on identifying occupancy breakpoints across signal phases, but in practice, such patterns are often obscured by noise and traffic heterogeneity. To address this, studies such as Cai et al. (2014); Wang et al. (2017), fuse probe vehicle data with detector measurements to estimate shockwave speeds, enabling the estimation of cycle-wise maximum and residual queues. An et al. (2018) proposed a hybrid approach: applying input-output models when breakpoints are unreliable, and using shockwave analysis otherwise. While these shockwave-based models offer a theoretical basis for queue evolution, they assume uniform traffic conditions, which rarely hold in urban settings.
Speed-based methods estimate the queue tail by analyzing speed patterns across different segments of a section over time, typically using floating car data (FCD). Some approaches detect the queue tail by identifying spatial speed drops alone from probe FCD (Dinh et al., 2014). As aFCD provides segment-level average speeds, Altıntaşı (2018) utilized it to generate synthetic probe trajectories using four-node quadratic interpolation and constructed iso-speed contours to estimate queue length. van Erp et al. (2018) combined aFCD with loop detector data to estimate vehicle accumulation. While such speed-based methods are intuitive and data-efficient, their accuracy heavily depends on the data quality and the speed thresholds used to identify congestion. Recently, Luo et al. (2022) eliminated the need for speed thresholds by employing multisource data with a back-propagation neural network optimized using the resilient propagation algorithm (RPROP-NN). In this approach, aFCD speeds and traffic light event data are first evaluated by a decision tree to determine whether a queue has formed. The loop detector occupancy rate is then assessed to verify whether the queue length exceeds its capacity. However, such neural network-based methods require specific training for each section and function as black-box models, limiting their transferability and interpretability.
Kalman filtering algorithms are based on state-space models, which describe system dynamics through a hidden state evolution equation and a measurement (observation) equation. The hidden state represents queue-length-related variables that evolve over time but are not directly observable. As is standard in state-space modeling, different data sources can be integrated into either the evolution model or the measurement model, providing flexibility. This framework is crucial for queue estimation because it incorporates noisy, indirect measurements from various sensors while explicitly modeling uncertainties in traffic dynamics and observations. The Kalman filter recursively estimates the hidden states by alternating between prediction—forecasting the system state and expected measurements—and update steps that refine estimates based on actual observations, providing a robust, real-time solution for tracking queue dynamics.
Several studies have adapted Kalman-based methods to queue length estimation. Abewickrema et al. (2023) applied an extended Kalman filter, modeling the cycle-wise maximum and residual queue lengths as the system state, which is assumed to remain static within each signal cycle. The measurement model is derived from shockwave theory to capture phase breakpoints. Hu et al. (2022) implemented two independent extended Kalman filters to separately track queue formation and discharge waves, using a measurement model that estimates shockwave location and speed from the hidden state. Wang et al. (2021) used a standard Kalman filter where the hidden state represents the number of queued vehicles. The evolution model is grounded in input-output theory, while the measurement model estimates queue length by accounting for the probe vehicle penetration rate. Similarly, Heshami and Kattan (2021) employed an evolution model based on input-output theory, but with a measurement model derived from a case-based reasoning algorithm.
Although traditional Kalman filters are interpretable and computationally efficient, they depend on full prior knowledge of system evolution dynamics and measurement models, which often do not match practical scenarios. For example, Noise covariances can be complex, unknown, and time-varying, which can degrade the performance of linear Kalman-based methods (Li et al., 2015). These limitations are particularly pronounced in queue length estimation, where traffic dynamics are inherently nonlinear, influenced by heterogeneous and time-varying factors (Ahmed et al., 2021; Chauhan et al., 2021) such as turning movements, and challenging to model accurately in partially observed traffic scenarios.
1.2 Research gap and contribution
There are three key limitations in current approaches. First, limitations in problem setup: most existing studies are developed and validated in well-controlled environments, often overlooking practical challenges such as partially observed traffic flows. In real-world urban settings, it is common for vehicles to enter or exit the study area from adjacent roads or unsignalized intersections without being monitored due to sensor installation constraints and failures (Federal Highway Administration, 2006; Shang et al., 2024). Second, data dependency: studies using loop detectors tend to rely on established models like input-output and shockwave theories, which impose strict requirements on data quality (e.g., accurate inflow/outflow counts and clear occupancy patterns) that are frequently unmet in real-world settings. On the other hand, aFCD—valued for its low cost, wide coverage, and privacy advantages—remains underutilized. As a result, current methods lack generalizability across different network sensor setups, with performance degrading significantly when applied to imperfect real world. Third, methodological constraints: many approaches depend on traditional Kalman filters and Extended Kalman filters, which impose linear or linearized models and known noise covariances. These assumptions are often unrealistic for dynamic traffic conditions at intersections, limiting the adaptability and scalability of such methods in diverse real-world applications.
To address these limitations, we propose Q-Net, a novel queue length estimation approach tailored to urban environments. First, Q-Net handles partially observed traffic flows, making it robust to incomplete monitoring coverage in real-world settings. Second, our approach leverages aggregated loop detector and aFCD data as inputs, both of which are widely available and privacy-friendly data sources. Third, to overcome the methodological constraints of traditional Kalman filters while preserving interpretability, we adopt an AI-augmented Kalman filtering approach that combines deep learning with state-space modeling. To achieve this, we adapt KalmanNet (Revach et al., 2022) to the queue length estimation problem. KalmanNet and its variants have gained attention in signal processing for their ability to learn complex dynamics while maintaining the recursive structure and transparency of classical filtering methods (Buchnik et al., 2023; Revach et al., 2023; Choi et al., 2023).
Our main contributions are as follows:
  1. 1.
    Interpretable state-space formulation: We develop a state-space model for queue length estimation that relies solely on loop detector data installed near the stop lines of adjacent intersections and aggregated Floating Car Data (aFCD) as inputs, both of which are low cost and widely available. While loop detectors can be flexibly positioned along the roadway, our method remains applicable by accounting for the time lag between vehicle detection and arrival at the stop line. The state-space model consists of an evolution model based on cumulative vehicle counts from loop detector data and a measurement model derived from aFCD speeds. It accommodates general traffic conditions including scenarios with partially observed traffic flows. This design enables effective fusion of data sources with differing spatial and temporal resolutions.
  2. 2.
    Data-driven solution: Q-Net addresses the trade-off between modeling capacity and interpretability of black-box neural networks. It integrates Kalman filtering principles with deep learning, following AI-augmented Kalman filtering techniques. Specifically, building on the tailored state-space formulation, Q-Net employs neural networks to calculate the Kalman gain and maintains physically meaningful internal states. Q-Net uses a measurement grouping strategy that makes model dimensions independent of road section length, which ensures transferability across different traffic scenarios—the model can be trained on one section and tested on another. Q-Net is highly data-efficient with only 952 parameters and trainable on 13-day datasets. A real-time learning mechanism is also incorporated for adaptive traffic control. To our knowledge, Q-Net represents the first application of AI-augmented Kalman filtering to queue length estimation.
  3. 3.
    Real-world validation: Q-Net has been validated on main urban road sections in Rotterdam, the Netherlands. Training on one road section and testing across different days yields over 60% RMSE improvement compared to baseline models, showing strong temporal transferability. The same trained model works directly on adjacent road sections without any finetuning. These adjacent sections present different traffic patterns due to their unique network configurations, yet Q-Net maintains reliable performance validated through quantitative evaluation and visual inspection.
The remainder of this paper is organized as follows: Section 2 introduces the fundamentals of the Kalman filter. Section 3 presents the problem setup and highlights key challenges. Section 4 describes how Q-Net models traffic dynamics using a state-space formulation, and Section 5 details its filtering procedure. Section 6 presents the implementation of Q-Net on major road sections in Rotterdam, analyzes the results, demonstrates its transferability, and showcases its real-time version for traffic control. Finally, Section 7 concludes the paper.
2 Background in Kalman filtering
State-space models describe how a system’s internal (latent) state and measurements evolve across time steps. Let $\mathbf{x}_{t}$ denote the latent state of the system at time $t$, $\mathbf{u}_{t}$ the control input, and $\mathbf{y}_{t}$ the observed measurement. The system is governed by:
(1) State evolution model that describes how the state evolves over time:
$\mathbf{x}_{t}=\mathbf{F}_{t}\mathbf{x}_{t-1}+\mathbf{G}_{t}\mathbf{u}_{t}+\mathbf{w}_{t},$ (1)
where $\mathbf{F}_{t}$ is the state transition matrix, $\mathbf{G}_{t}$ is the control input matrix, and $\mathbf{w}_{t}\sim\mathcal{N}(\mathbf{0},\mathbf{Q}_{t})$ is zero-mean Gaussian process noise.
(2) Measurement model that relates the observed measurement to the current state:
$\mathbf{y}_{t}=\mathbf{H}_{t}\mathbf{x}_{t}+\mathbf{v}_{t},$ (2)
where $\mathbf{H}_{t}$ is the observation matrix and $\mathbf{v}_{t}\sim\mathcal{N}(\mathbf{0},\mathbf{R}_{t})$ is zero-mean Gaussian measurement noise. The noise $\mathbf{w}_{t}$ and $\mathbf{v}_{t}$ are modeled to be mutually independent and also independent of both the states and the measurements. The initial state $\mathbf{x}_{0}$ is modeled to be Gaussian, with known mean $\hat{\mathbf{x}}_{0}$ and covariance $\hat{\boldsymbol{\Sigma}}_{0}$.
The Kalman filter recursively estimates the a posteriori distribution of the latent state $\mathbf{x}_{t}$ given all past control inputs and measurements up to time $t$. Let $\mathbf{\Sigma}_{t}$ and $\mathbf{S}_{t}$ denote the covariance matrices of the state and the measurement, respectively. At each time step, the algorithm proceeds in two stages: prediction and update.
Prediction. The filter first computes a priori estimates of the state $\hat{\mathbf{x}}_{t|t-1}$ and its uncertainty $\boldsymbol{\Sigma}_{t|t-1}$ based on the previous a posteriori state estimate $\hat{\mathbf{x}}_{t-1|t-1}$ and uncertainty estimate $\boldsymbol{\Sigma}_{t-1|t-1}$:
$\hat{\mathbf{x}}_{t|t-1}$ $=\mathbf{F}_{t}\hat{\mathbf{x}}_{t-1|t-1}+\mathbf{G}_{t}\mathbf{u}_{t},$ (3)
$\boldsymbol{\Sigma}_{t|t-1}$ $=\mathbf{F}_{t}\boldsymbol{\Sigma}_{t-1|t-1}\mathbf{F}_{t}^{\top}+\mathbf{Q}_{t}.$ (4)
Throughout this paper, $(\cdot)^{\top}$ denotes the matrix transpose operation. The predicted measurement is:
$\hat{\mathbf{y}}_{t|t-1}$ $=\mathbf{H}_{t}\hat{\mathbf{x}}_{t|t-1},$ (5)
$\mathbf{S}_{t|t-1}$ $=\mathbf{H}_{t}\boldsymbol{\Sigma}_{t|t-1}\mathbf{H}_{t}^{\top}+\mathbf{R}_{t}.$ (6)
Update. Using the new measurement $\mathbf{y}_{t}$, the filter updates the state estimate:
$\hat{\mathbf{x}}_{t|t}$ $=\hat{\mathbf{x}}_{t|t-1}+\mathbf{\mathcal{K}}_{t}\Delta\mathbf{y}_{t},$ (7)
$\mathbf{\Sigma}_{t|t}$ $=\mathbf{\Sigma}_{t|t-1}-\mathbf{\mathcal{K}}_{t}\mathbf{S}_{t|t-1}\mathbf{\mathcal{K}}_{t}^{T}.$ (8)
The innovation is given by $\Delta\mathbf{y}_{t}=\mathbf{y}_{t}-\hat{\mathbf{y}}_{t|t-1}$, representing the prediction error between the observed measurement $y_{t}$ and the one-step-ahead prediction $\hat{y}_{t|t-1}$. The Kalman gain $\mathbf{\mathcal{K}}_{t}$ is defined as $\mathbf{\mathcal{K}}_{t}=\mathbf{\Sigma}_{t|t-1}\mathbf{H}^{T}\mathbf{S}_{t|t-1}^{-1}$. It is chosen to minimize the mean squared error (MSE) of the state estimate. The classical Kalman filter thus takes the posteriori first- and second-order moments from the previous time step and produces those for the current step, enabling online, recursive state estimation. Another widely used extension is the Extended Kalman Filter (EKF), which handles nonlinear systems by locally linearizing the state and measurement models via their Jacobian matrices.
3 Problem setup and challenges
A lane refers to a portion of the roadway designated for a single line of vehicles, which serves to organize traffic, guide drivers, and reduce conflicts. A section is a group of lanes that share the same direction of travel. As illustrated in Figure 1, a single lane may accommodate multiple movement directions, such as through and right-turn traffic. A section typically contains multiple lanes, which may diverge into different movement directions at the downstream intersection.
Figure 1: Layout of a signalized intersection, including the definitions of a lane, section, and ground truth queue length. Data are collected from inflow/outflow detectors and aFCD segments. The layout highlights two common sources of partially observed traffic flow: (i) lanes with combined directions at the upstream intersection of the study section, where part of the traffic flow does not proceed to the downstream area, (ii) unobserved traffic flow leaving or entering the study section. The leaving flow can be caused by infrastructure splits (e.g., elevated roads or tunnels) or vehicles entering neighboring buildings, while the entering flow consists of minor car movements from the surrounding environment.
Our problem of interest is to estimate queue lengths at signalized intersections using indirect data sources that do not explicitly measure queues. In urban road traffic networks, direct queue measurement is often impractical or costly, yet accurate queue length estimation is essential for traffic management. We focus on section-wise queue length estimation, where the queue length represents the maximum queuing distance across all lanes within a section.
The problem involves fusing heterogeneous data sources with different characteristics. Loop detectors installed near stop lines record vehicle passages. These passages are aggregated into 10-second vehicle counts. The aFCD divides each road section into segments and provides average speed data for each segment at 1-minute intervals. Radar sensors capture ground truth queue lengths at 10-second intervals. A summary is presented in Table 1.
This presents several key challenges from both road network structure and data characteristics:
  1. 1.
    Partially observed traffic flow. Figure 1 illustrates scenarios leading to unobserved vehicle movements: (i) vehicles not entering downstream areas due to multi-directional upstream lanes, and (ii) vehicles exiting or entering the study area via neighboring environments that bypass the monitored network. Entering vehicles typically form minor flows, while exiting vehicles contribute to secondary flows. These unmonitored movements introduce uncertainty into queue length estimation.
  2. 2.
    Spatiotemporal data inconsistency. The data sources exhibit significant heterogeneity in both spatial granularity and temporal resolution, as shown in Table 1. This inconsistency requires models capable of reconciling spatial and temporal mismatches for effective data fusion.
Our goal is to develop a data-driven framework that effectively combines these multi-resolution, multi-source data to produce accurate queue length estimates. The following section presents our proposed approach to address these challenges, highlighting methodological innovations and practical relevance for real-world applications.
Table 1: Overview of data sources and their characteristics
Input Data Ground Truth
Data Source Loop Detector aFCD Radar
Traffic Variable Vehicle Count Average Speed Queue Length
Spatial Coverage Lane-level at stop lines (upstream and downstream) Section segments Section-level
Temporal Resolution 10 seconds 60 seconds 10 seconds
Purpose Model Input Model Input Evaluation
4 Q-Net state-space formulation
Let $x_{t}\in\mathbb{R}$ be the section-level queue length at time step $t$ as the state variable, let $u_{t}\in\mathbb{R}$ be the queue change as the control input, and let $\mathbf{y}_{t}=[y_{t}^{1},y_{t}^{2},\dots,y_{t}^{N}]^{\top}\in\mathbb{R}^{N}$ be the segment-wise aFCD speeds as measurements, where $N$ is the number of section segments. Our state-space model has the form:
$x_{t}$ $=x_{t-1}+u_{t}+w_{t},$ (9)
$\mathbf{y}_{t}$ $=\mathbf{h}(x_{t})+\mathbf{v}_{t}.$ (10)
This formulation seamlessly combines spatiotemporally inconsistent data sources within the Q-Net framework. The evolution equation (9) expresses the queue length $x_{t}$ at time $t$ as the sum of the previous queue length $x_{t-1}$, the queue change $u_{t}$, and the process noise $w_{t}$, while the measurement equation (10) establishes the relationship between aFCD speeds $\mathbf{y}_{t}$ and queue length $x_{t}$ through a nonlinear mapping $\mathbf{h}(\cdot)$.
The remaining model components are defined as follows: $w_{t}\sim\mathcal{N}(0,Q_{t})$ is Gaussian process noise with covariance $Q_{t}$, and $\mathbf{v}_{t}\sim\mathcal{N}(\mathbf{0},\mathbf{R}_{t})$ is Gaussian measurement noise with covariance $\mathbf{R}_{t}$. The control input $u_{t}$ (queue change) is derived from cumulative loop detector counts (see Subsection 4.1), and $\mathbf{h}(\cdot)$ is a nonlinear mapping function detailed in Subsection 4.2.
We do not treat noise covariances as fixed or known; instead, $Q_{t}$ and $\mathbf{R}_{t}$ are treated as time-varying and learnable parameters. The covariance of state variable and measurements are denoted as $\Sigma_{t}$ and $\mathbf{S}_{t}$. A summary of the notations used throughout this paper is provided in Table 2.
Table 2: Overview of notation, descriptions, and units.
Notation Description Unit
$t$ Discrete time index
$x_{t}$ State variable at time $t$: section-level queue length m
$u_{t}$ Control input at time $t$: estimated queue change m
$\mathbf{y}_{t}$ Measurement vector at time $t$ (aFCD speeds) m/s
$y_{t}^{i}$ Speed measurement for $i$-th section segment at time $t$ m/s
NN Number of aFCD segments within the section
$w_{t},Q_{t}$ Process noise and its covariance matrix m, m²
$\mathbf{v}_{t},\mathbf{R}_{t}$ Measurement noise and its covariance matrix m/s, (m/s)²
$\mathbf{h}$ Measurement model function
$\Sigma_{t}$ Second-order statistical moment of state variable $x_{t}$
$\mathbf{S}_{t}$ Second-order statistical moment of measurements $\mathbf{y}_{t}$ (m/s)²
$\lambda_{c}$ Unobserved traffic flow rate vehicles/s
$m$ Number of lanes in the section
$L$ Length of the road section m
$A_{t}$ Cumulative inflow (vehicle arrivals) count from upstream vehicles
$D_{t}$ Cumulative outflow (vehicle departures) count from the section vehicles
$k_{j},k_{f}$ Jam density and free-flow density vehicles/m
$v_{j},v_{f}$ Jam speed and free-flow speed m/s
$l^{i},r^{i}$ Start and end positions of segment $i$ measured from the stop line m
$Q_{\max}$ Maximum queue length for the section m
$\hat{x}_{t|t}$ Posteriori estimate of $x_{t}$ given mesurements up to time $t$ m
$\hat{y}_{t|t-1}$ Predicted measurement at time $t$ based on state at time t1t-1 m/s
$\mathcal{K}_{t}(\theta)$ Kalman gain parameterized by $\theta$
4.1 Control queue change input
In this subsection, we explain how to derive the control input $\mathbf{u}_{t}$ in Equation (9). The approach involves recovering queue length using only cumulative vehicle counts and utilizing the temporal differences as queue changes. Since we are primarily interested in the control input (i.e., queue change), our goal is not to reproduce absolute queue length values, but rather to capture the temporal dynamics of queue buildup and dissipation. Let $q_{t}$ denote the reconstructed queue length at time $t$. The control input is derived as:
$u_{t}=\mathcal{D}(\mathcal{F}(\mathcal{C}(q_{t}))),$ (11)
where $\mathcal{C}$ is the operation to account for unobserved traffic flow, $\mathcal{F}$ is the Fourier band-pass filtering operation, and $\mathcal{D}$ is the temporal difference operation.
Queue length recovery considering unobserved traffic. Let $\lambda_{c}$ be the unobserved net traffic flow rate, defined as the net number of vehicles leaving the system per unit time without passing through the designated outflow point, partially offset by unobserved minor traffic inflow. We model the section with length $L$ and $m$ lanes with equally formed queues, where the queued portion operates at jam density $k_{j}$ and speed $v_{j}$, while the remaining portion operates at free-flow density $k_{f}$ and speed $v_{f}$. Let $A_{t}$ and $D_{t}$ respectively be cumulative vehicle inflow and outflow up to time $t$. Applying macroscopic vehicle conservation:
$mq_{t}k_{j}+m(L-q_{t})k_{f}=A_{t}-D_{t}-\lambda_{c}t.$ (12)
The lefthand side represents the total amount of vehicles in the section: $mq_{t}k_{j}$ in the queued region and $m(L-q_{t})k_{f}$ in the free-flow region. The righthand side accounts for net vehicle accumulation adjusted for unobserved flows. Equation (12) yields:
$q_{t}=\frac{A_{t}-D_{t}-mLk_{f}-\lambda_{c}t}{m(k_{j}-k_{f})}.$ (13)
Since $A_{t}-D_{t}$ can be directly obtained from loop detector counts, the only remaining unknown that governs the temporal evolution of $q_{t}$ is the unobserved net flow rate $\lambda_{c}$
We estimate $\lambda_{c}$ using boundary conditions where queue length is approximately zero at the beginning and end of each analysis period (e.g., start and end of day). Once $\lambda_{c}$ is identified, rather than fixing the values for $k_{j}$ and $k_{f}$, we apply an affine transformation to $A_{t}-D_{t}-\lambda_{c}t$ through shifting and scaling so that the adjusted signal spans the empirically observed queue length range $0,Q_{\max}]$, where $Q_{\max}$ is the section-specific maximum. This approach avoids assumptions about jam and free-flow densities, allowing the model to learn these parameters directly from daily vehicle count patterns. The reconstructed queue length is shown in Figure 2(a).
(a) Orange line: Raw net vehicle accumulation $(A_{t}-D_{t})$. Blue line: Reconstructed queue length by removing the unobserved traffic flow from the net flow and applying an affine transformation to align with the observed queue range.
(b) Orange line: Reconstructed queue length using a Fourier filter that suppresses low-frequency drift and high-frequency noise, producing a smoothed queue profile suitable for control input $u_{t}$. Blue line: Ground truth queue length.
Figure 2: Queue estimation from net vehicle accumulation under unobserved traffic flow conditions for section N1-IN in Rotterdam on 2023-11-10.
Control input derivation from reconstructed queue length. From Figure 2(a), the previous reconstructed queue length exhibits a slow-varying trend over time. This low-frequency drift is likely due to the fact that the unobserved traffic flow rate $\lambda_{c}$ is not strictly constant over time—e.g., due to changes in turning ratios at different periods during the day. Additionally, high-frequency noise may arise from detector errors and aggregation mismatches. We apply a Fourier-based bandpass filter to suppress both low-frequency drift and high-frequency noise, emphasizing medium-frequency variations most representative of actual queue dynamics. Figure 2(b) shows the filtered reconstructed queue length. Since our goal is to approximate queue changes (the control input) rather than exact magnitudes, even if the average estimated queue during the afternoon peak appears similar to that in the morning, this is of limited importance because what matters are the relative changes.
4.2 Measurement function
We use our two-regime traffic behavior stated in Subsection 4.1: when a portion of a section is queued, vehicles move at a congested speed $v_{j}$; when it is not queued, they travel at a free-flow speed $v_{f}$. Let segment $i$ extend from position $l^{i}$ to $r^{i}$ (with both measured as distances from the stop line and $r^{i}>l^{i}$). At time $t$, the expected speed observation $y_{t}^{i}$ for this segment is computed as:
[…]
The equation shows that fully unqueued or queued segments have expected speed observation of $v_{f}$ or $v_{j}$. For partially queued segments, where the queue boundary lies within the segment, the average speed is calculated as segment length divided by total travel time, which is the sum of time in the queued portion (traveled at $v_{j}$) and unqueued portion (traveled at $v_{f}$). Applying this formulation to all segments, we build a mapping function from queue length $x_{t}$ to the measurement vector $\mathbf{{y}}_{t}=[{y}^{1}_{t},{y}^{2}_{t},\dots,{y}^{N}_{t}]^{\top}$.
The characteristic speeds, $v_{j}$ and $v_{f}$, are empirically estimated by picking the two dominant modes from the observed aFCD speed distribution through histogram analysis.
Figure 3: At time $t$, the Q-Net filtering framework takes three inputs: the previous a posteriori state estimate $\hat{x}_{t-1|t-1}$, cumulative vehicle counts $A_{t},D_{t}$, and aFCD speed measurements $\mathbf{y}_{t}$, and produces one output: the current a posteriori state estimate $\hat{x}_{t|t}$. Predict: $A_{t}$ and $D_{t}$ are used to derive the control input (queue change) ${u}_{t}$, which predicts the a priori queue length $x_{t|t-1}$ and a priori speed measurements $\mathbf{\hat{y}_{t|t-1}}$ through the measurement model. Update: The predicted a priori speeds $\mathbf{\hat{y}_{t|t-1}}$ are compared with actual aFCD measurements $\mathbf{y}_{t}$, and the adapted KalmanNet calculates the Kalman gain $\mathcal{K}_{t}$ to update the a posteriori queue estimate $\hat{{x}}_{t|t}$.
5 Queue-length Estimation framework
We now detail the two-step filtering procedure of Q-Net: predict and update. As illustrated in Figure 3 and in Algorithm 1, Q-Net operates similarly to traditional Kalman filters through a predict-update loop.
Algorithm 1 Queue-length Estimation Algorithm
Input: Initial state $\hat{x}_{0|0}$, control inputs $\{u_{t}\}$, measurements $\{\mathbf{y}_{t}\}$, maximum queue length $Q_{\max}$, parameters $\theta$
Output: Queue length estimates $\{\hat{x}_{t|t}\}$
For $t=1,2,\ldots,T$:
  1. 1.
    Predict:
    • $\hat{x}_{t|t-1}=\min(\max(\hat{x}_{t|t-1}+u_{t},0),Q_{\max})$
    • $\hat{\mathbf{y}}_{t|t-1}=\mathbf{h}(\hat{x}_{t|t-1})$
  2. 2.
    Update:
    • $\mathcal{K}_{t}=\text{NeuralNetwork}(\theta,\text{context}_{t})$
    • $\hat{x}_{t|t}=\hat{x}_{t|t-1}+\mathcal{K}_{t}(\mathbf{y}_{t}-\hat{\mathbf{y}}_{t|t-1})$
    • $\hat{x}_{t|t}=\min(\max(\hat{x}_{t|t},0),Q_{\max})$
Predict. The predict step estimates the a priori state variable (i.e., the queue length) $\hat{x}_{t|t-1}$ using the previous a posteriori state estimate $\hat{x}_{t-1|t-1}$ and the control input $u_{t}$ (i.e., the queue change). At the same time, it predicts the expected measurement $\hat{\mathbf{y}}_{t|t-1}$ based on the current a priori state through the measurement model $\mathbf{h}(\cdot)$. The predict step is given by:
$\hat{x}_{t|t-1}=$ $\min(\max(\hat{x}_{t-1|t-1}+u_{t},0),Q_{\max}),$ (15)
$\hat{\mathbf{y}}_{t|t-1}$ $=\mathbf{h}(\hat{x}_{t|t-1}).$ (16)
A clipping operation is applied to ensure that the state variable remains within the range $0,Q_{\max}]$.
Update. As for the update step, the system state is updated using the innovation (i.e., the difference between the predicted and observed measurements) scaled by the Kalman gain $\mathcal{K}_{t}$. Classical Kalman filtering imposes linear evolution and measurement models, along with fixed Gaussian noise. Under these assumptions, the Kalman gain can be computed analytically from the second-order statistics of the state and observation variables. However, in queue length estimation, the state-space model is nonlinear (see Section 4), and the noise characteristics can vary significantly over time. For instance, as shown in Figure 2, the process noise during the afternoon differs significantly from that in the morning. To address these challenges, we learn the Kalman gain $\mathcal{K}_{t}(\theta)$ parameterized by $\theta$ directly from the data following the KalmanNet principle. After receiving the true measurement $\mathbf{y}_{t}$, the system state is updated by:
$\hat{x}_{t|t}=\min(\max(\hat{x}_{t|t-1}+\mathcal{K}_{t}(\theta)(\mathbf{y}_{t}-\mathbf{\hat{y}}_{t|t-1}),0),Q_{\max}).$ (17)
The min-max operation implements state clamping compared to the classical Kalman filter update (see Equation (7)). While this modification sacrifices mathematical optimality, it ensures state estimates remain within physically meaningful bounds $0,Q_{\max}]$. This approach is widely accepted in engineering practice due to its simplicity and computational efficiency (Simon, 2010; Prakash et al., 2014). Since the true measurements (aFCD) have a coarser temporal resolution (60s) compared to the state variable (queue length estimated every 10s), we use the same speed measurement for all time steps within the corresponding one-minute interval.
Within this framework, Q-Net integrates data sources of different resolutions: loop detector data, available at high temporal resolution, is used in the predict step to model queue evolution, while coarse aFCD speed measurements are used to correct predictions in the update step. This design ensures that the overall trend of queue dynamics is shaped by speed observations, while finer temporal variations are captured through detailed loop detector data.
The detailed computation of $\mathcal{K}_{t}(\theta)$ is detailed in Subsection 5.1, and the adaptations to enforce spatial transferability are described in Subsection 5.2.
5.1 Learned Kalman gain
Equation (17) highlights a key departure of Q-Net from traditional Kalman filters: instead of computing the Kalman gain analytically from model assumptions, Q-Net learns it directly from data. Drawing inspiration from the traditional Kalman filter, where the Kalman gain depends on the statistics of the state estimates and measurements, four features are designed in Revach et al. (2022) to capture the unknown statistical relationships between the state-space model and the Kalman gain:
  • The measurement difference $\Delta\tilde{\mathbf{y}}_{t}=\mathbf{y}_{t}-\mathbf{y}_{t-1}$;
  • The innovation difference $\Delta{\mathbf{y}}_{t}=\mathbf{y}_{t}-\mathbf{\hat{y}}_{t|t-1}$;
  • The forward evolution difference $\Delta\tilde{x}_{t}=\hat{x}_{t|t}-\hat{x}_{t-1|t-1}$;
  • The forward update difference $\Delta\hat{x}_{t}=\hat{x}_{t|t}-\hat{x}_{t|t-1}$.
The forward evolution difference $\Delta\tilde{x}_{t}$ represents the change between two consecutive a posteriori state estimates, while the forward update difference $\Delta\hat{x}_{t}$ captures the correction from the a priori to the a posteriori estimate at the same time step. However, since the a posteriori estimate $\hat{x}_{t|t}$ is not available before the Kalman gain is computed, we use the previous values $\Delta\tilde{x}_{t-1}$ and $\Delta\hat{x}_{t-1}$ during inference. In this setup, measurement difference $\Delta\tilde{\mathbf{y}}_{t}$ and forward evolution difference $\Delta\tilde{x}_{t-1}$ capture the underlying state evolution dynamics, whereas the innovation difference $\Delta{\mathbf{y}}_{t}$ and the forward update difference $\Delta\hat{x}_{t-1}$ reflect the uncertainty associated with the current state estimate.
Figure 4: KalmanNet architecture for dynamic Kalman gain estimation. The top panel illustrates the overall KalmanNet pipeline, which takes in four temporal input features (measurement difference $\Delta\tilde{\mathbf{y}}_{t}$, innovation difference $\Delta{\mathbf{y}}_{t}$, forward evolution difference $\Delta\tilde{x}_{t-1}$, forward update difference $\Delta\hat{x}_{t-1}$) to estimate the Kalman gain. Intermediate representations ($\hat{Q}_{t},\hat{\Sigma}_{t},\hat{\mathbf{S}}_{t}$ and $\hat{\Sigma}_{t}^{h}$) are computed through sequential submodules. The lower panels detail the submodules a, b, c, d, and e.
As shown in Figure 4, KalmanNet can be globally divided into five submodules:
  1. a.
    Process noise covariance module: This module estimates the process noise covariance matrix $\hat{Q}_{t}$ based on the forward update difference $\Delta\hat{x}_{t-1}$. The design is motivated by Equations (1) and (3), which suggest that evolution noise influences the predicted state estimate. The estimation follows:
    $\hat{Q}_{t}=\text{GRU}(\text{FC}(\Delta\hat{x}_{t-1})).$ (18)
    The fully connected (FC) layers transform the input into a higher-dimensional feature representation. The gated recurrent unit (GRU) maintains a hidden state that is updated at each time step through its gating mechanism, enabling $\hat{Q}_{t}$ to be estimated based on the complete history of forward update differences up to the current time rather than using fixed noise models.
  2. b.
    State covariance module: This module estimates the second-order moments of the model state, denoted as $\hat{\Sigma}_{t}$. According to Equation (4), the a posteriori state covariance depends on the process noise covariance $\hat{Q}_{t}$. Additionally, the forward evolution difference $\Delta\tilde{x}_{t-1}$ is included to capture the queue length evolution over time and provides system dynamics information that is not captured by process noise alone. The estimation is performed as follows:
    $\hat{\Sigma}_{t},\hat{\Sigma}_{t}^{h}=\text{GRU}(\text{Concat}[\hat{Q}_{t},\text{FC}(\Delta\tilde{x}_{t-1})]),$ (19)
    where $\hat{\Sigma}_{t}^{h}$ denotes the hidden state of the GRU, later refined in the hidden state update module, and Concat denotes concatenation. The GRU’s gating mechanism selectively retains relevant historical queue length covariance patterns while discarding outdated information.
  3. c.
    Measurement covariance module: This module estimates the measurement covariance matrix $\hat{\mathbf{S}}_{t}$. From Equation (6), $\hat{\mathbf{S}}_{t}$ is related to the innovation difference $\Delta\mathbf{y}_{t}$, which reflects speed prediction errors, and the estimated state covariance $\hat{\Sigma}_{t}$, which indicates queue length uncertainty. We also include the measurement difference $\Delta\tilde{\mathbf{y}}_{t}$ as input, as it captures temporal variations in raw aFCD speed measurement:
    $\hat{\mathbf{S}}_{t}=\text{GRU}(\text{Concat}[\text{FC}(\hat{\Sigma}_{t}),\text{FC}(\text{Concat}[\Delta\tilde{\mathbf{y}}_{t},\Delta\mathbf{y}_{t}])]).$ (20)
    The GRU uses a reset gate to determine how much of the past measurement covariance pattern remains valid, and an update gate to control how much the current uncertainty from speed prediction errors, queue length uncertainty, and speed variations should influence the measurement covariance estimate.
  4. d.
    Kalman gain module: This module computes the Kalman gain $\mathcal{K}_{t}$ based on the queue length and speed measurement covariances, consistent with the input of Kalman gain calculation in traditional Kalman filters:
    $\mathcal{K}_{t}=\text{FC}(\text{Concat}[\hat{\Sigma}_{t},\hat{\mathbf{S}}_{t}]).$ (21)
    Equations (3)-(8) show the recursive nature of Kalman filtering, where each Kalman gain updates the state estimate, which subsequently influences the next gain computation. This recursive nature of Kalman gain calculation is reflected in the design of the state covariance module (for $\hat{\Sigma}_{t}$) and measurement covariance module (for $\hat{\mathbf{S}}_{t}$), both of which use GRUs to capture temporal dependencies in the filtering process.
  5. e.
    Hidden state covariance update module: This module updates the internal hidden state $\hat{\Sigma}^{h}_{t}$ to retain temporal context across time steps, which is reflected by Equation (8):
    $\hat{\Sigma}^{h}_{t}=\text{FC}(\text{Concat}[\hat{\Sigma}_{t},\text{FC}(\text{Concat}[\mathcal{K}_{t},\hat{\mathbf{S}}_{t}])]).$ (22)
In this way, KalmanNet dynamically tracks the underlying statistical moments associated with state evolution and measurements. The Kalman gain is computed adaptively, leveraging both memory of past information and current predicted state and measurement data. For additional implementation details, please refer to Revach et al. (2022).
5.2 Adaptation for spatial transferability
Recall that a segment is the spatial unit over which aFCD aggregates speed measurements. While KalmanNet offers flexibility in nonlinear Kalman filtering without prior noise assumptions, it requires fixed-dimensional measurement vectors corresponding to the number of segments. Since obtaining ground truth data for training can be difficult or costly, transferring models across sections with varying segment counts is essential, necessitating spatial transferability adaptation.
To address this, we impose that the measurement input to KalmanNet remains fixed-dimensional regardless of the number of road segments, enabling Q-Net to be trained on one section and applied to others with spatial transferability. Specifically, we assume that the covariance entries for each segment depend only on the segment itself and its immediate neighbors (previous and next segments), which is justified by the short 10-second filtering interval (ground truth frequency). As shown in Figure 5, we adopt a local measurement grouping strategy, where each measurement group for segment $1<i<N$ is defined as:
  • The grouped measurement difference $\Delta\breve{\mathbf{y}}^{i}_{t}=[{y}_{t}^{i-1}-y_{t-1}^{i-1},{y}_{t}^{i}-y_{t-1}^{i},{y}_{t}^{i+1}-y_{t-1}^{i+1}]^{\top}$;
  • The grouped innovation difference $\Delta{\mathbf{\bar{y}}}^{i}_{t}=[{y}_{t}^{i-1}-{\hat{y}}_{t|t-1}^{i-1},{y}_{t}^{i}-{\hat{y}}_{t|t-1}^{i},{y}_{t}^{i+1}-{\hat{y}}_{t|t-1}^{i+1}]^{\top}$.
We exclude boundary segments ($i=1$ and $i=N$) from the update process, as queue dynamics are primarily reflected by intermediate segments. The first segment near the stop line can be erroneous due to localized queue dissipation patterns, while the last segment provides limited information since intermediate segments repeatedly confirm the presence of a long queue when it extends to the section boundary.
Figure 5: Local measurement grouping strategy for transferable Kalman gain estimation. aFCD speed measurements are grouped into fixed-size vectors consisting of the previous, current, and next segments. Each vector is processed by KalmanNet to obtain a local Kalman gain, and the final state update is computed as the sum of all local updates. This approach enables model transfer across road sections with different numbers of segments.
These grouped features are input to KalmanNet, which produces a group-specific Kalman gain $\mathcal{K}^{i}_{t}$. Note that each group is not expected to share the same covariance matrix; instead, the model learns a parameterized function that is shared across groups, enabling adaptive estimation of the covariance structure for each local neighborhood. The state estimate is then updated by aggregating the additive contributions of all groups, preserving the structure of the classical Kalman filter:
$\hat{x}_{t|t}=\hat{x}_{t|t-1}+\sum_{i=2}^{N-1}\mathcal{K}^{i}_{t}\Delta\bar{\mathbf{y}}^{i}_{t}$ (23)
Although individual measurements may appear in multiple overlapping groups, KalmanNet learns appropriate group-specific gains that implicitly balance their influence. This allows the network to determine how much each local group should affect the global state estimate. Moreover, this grouping strategy enables KalmanNet to generalize across road sections with varying segmentations. By focusing on the current segment and its adjacent segments, the model captures local neighborhood covariances, enhancing both robustness and transferability.
6 Numerical results
In this section, we validate our method Q-Net using a real-world case study from Rotterdam, the Netherlands. The dataset spans 18 days (from 6:00 AM to 8:00PM), from 2023-11-10 to 2023-11-27, comprising 12 weekdays and 6 weekend days. We randomly split the data into training, validation, and test sets while ensuring a balanced representation of weekdays and weekends. Specifically, 2023-11-11 (Saturday) and 2023-11-20 (Monday) are used for validation, while 2023-11-14 (Tuesday), 2023-11-18 (Saturday), and 2023-11-22 (Wednesday) are designated as test days.
We present the experimental setup in Subsection 6.1, evaluate model performance against baselines in Subsection 6.2, assess transferability in Subsection 6.3, demonstrate real-time applicability in Subsection 6.4, and analyze component contributions through ablation studies in Subsection 6.5.
6.1 Experiment setup
Q-Net is trained end-to-end in a supervised fashion and optimized using the Adam optimizer. When aFCD data is missing, the missing value is imputed by the previous value of the same segment. The loss function is defined as the Root Mean Square Error (RMSE) between the estimated queue length $\hat{x}_{t|t}$ and the ground truth queue length $x_{t}$ measured by radar data. Given a sequence of $T$ time steps, the loss is computed as:
$\mathcal{L}=\sqrt{\frac{\sum_{t=0}^{T-1}(x_{t}-\hat{x}_{t|t}^{2})^{2}}{T}}$ (24)
During training, the clipping functions in Equations (15) and (17) are disabled to ensure uninterrupted gradient backpropagation. These functions are applied only during validation and testing to enforce realistic queue length constraints.
In this study, each time series spans from 6:00 AM to 8:00 PM. With a time resolution of 10 seconds, this results in 5,040 time steps per sequence. Due to the long sequence length, accumulated prediction errors during the early stages of training may hinder effective learning. To mitigate this, we divide the sequence into shorter windows of 10 minutes for neural network training and model weight updates. The system state at the beginning of each window is initialized to zero if it is the first window of the day; otherwise, it is initialized using the a posteriori state estimate from the preceding window.
Q-Net is computationally efficient, comprising only 952 trainable parameters and demonstrating fast training performance. When trained on an Apple M2 Pro chip using 13 days of data, each epoch takes approximately 40 seconds. The cutoff frequencies of the Fourier-based bandpass filter can be empirically determined by visual inspection of the filtered data. The KalmanNet implementation follows that of Revach et al. (2022) [^1] https://github.com/KalmanNet/KalmanNet_TSP, with the learning rate set to 0.001.
We compare our model against baseline approaches. First, we review prior work using similar datasets, namely loop detector counts and/or aFCD speed data. For example, Luo et al. (2022) combines aFCD with loop detector data but also requires traffic light event data. Models relying solely on loop detector data are typically designed for cycle-wise maximum queue length estimation (Liu et al., 2009; Abewickrema et al., 2023; Sengupta et al., 2021; An et al., 2018), or assume detectors placed at the entrance of the section with fully observed and conserved traffic flow (Heshami and Kattan, 2021; Lee et al., 2019; Yang et al., 2019; Karnati et al., 2021). These assumptions are not compatible with our study context, which includes a large unobserved traffic flow and limited loops. As a result, we select two simple yet representative methods that rely only on aFCD as our baselines:
  • OSD (Altıntaşı, 2018), which detects the end of the queue based on observed speed drops where the upstream speed is higher than 16 km/h and downstream speed is lower than 16 km/h.
  • ISC (Dinh et al., 2014), which constructs synthetic trajectories from aFCD and estimates the queue using iso-speed contours. Speed thresholds are selected to be 14 km/h, 16 km/h, 18 km/h, 20 km/h, and 22 km/h.
6.2 Performance
Figure 6: Satellite view of N1-IN. The downstream area is divided into the study section (highlighted in dark red) and a grade-separated interchange. This interchange links to an open bridge, which serves as a critical entry point to the Rotterdam urban area. Therefore, the unobserved traffic flow in this region is significant.
Figure 7: Box plots of absolute estimation errors for baseline models (OSD, ISC) and Q-Net across all test days for section N1-IN. All p-values between the error distributions of baseline methods and our approach are less than $10^{-5}$, indicating statistically significant differences.
Rotterdam is the second-largest city in the Netherlands. To evaluate our method, we first filter out road segments lacking sufficient data availability, either due to contractual restrictions with the municipality or the presence of significant missing values. Among the remaining candidates, we select N1-IN as it is one of the major urban road sections in the Rotterdam traffic network. Evaluating our model on this section is of practical importance because it directly connects to the tunnel in Rotterdam, a critical point of interest for the municipality’s traffic management efforts. A satellite view of the section is shown in Figure 6. After vehicles exit the upstream intersection, they can either enter the ring interchange or continue through N1-IN, leading to significant unobserved traffic flow exiting the study section.
Table 3: Performance comparison between the proposed model and baseline models in terms of RMSE, MAE, and MAPE metrics.
All day Morning peak Afternoon peak
RMSE MAE MAPE RMSE MAE MAPE RMSE MAE MAPE
OSD 214.62 130.57 104.25 18.74 50.15 130.32 247.09 308.21 69.96
ISC 188.99 127.29 95.78 34.20 68.41 130.38 245.97 269.03 59.74
Q-Net (ours) 71.65 49.51 48.60 5.29 36.94 79.43 26.79 68.04 17.91
Improvement 62.09% 61.10% 49.26% 71.77% 26.34% 39.05% 89.11% 74.71% 70.02%
In Table 3, we evaluate model performance on the test days using three metrics: Root Mean Square Error (RMSE), Mean Absolute Error (MAE), and Mean Absolute Percentage Error (MAPE). For MAPE, we consider only instances where the ground truth queue length exceeds 10 meters to avoid skewed results from disproportionately large percentage errors on near-zero values. Considering that traffic dynamics vary substantially across time periods, metrics are reported for the entire day, the morning peak, and the afternoon peak. The morning and afternoon peak hours are defined according to the guidelines from the Dutch Ministry of Infrastructure and Water Management (Mobiliteitsscan, ).
Our results show that the proposed Q-Net consistently outperforms the baseline approaches across all metrics and time periods. We observe that MAPE is generally higher during the morning peak hours. Upon reviewing the ground truth data, we found that substantial queues during the morning peak occurred only on 2023-11-14, whereas the other test days exhibited much shorter queues. As MAPE is sensitive to low ground truth values, shorter queues naturally result in higher percentage errors.
(a) The aFCD.
(b) Q-Net vs OSD.
(c) Q-Net vs ISC.
Figure 8: The aFCD visualization and queue length estimation comparison among baselines (OSD and ISC), and our proposed method for section N1-IN on 2023-11-14.
Figure 7 shows the box plot of absolute estimation errors for each method across all test days. Q-Net demonstrates significantly lower mean errors and reduced variance. Notably, all methods show smaller errors on 2023-11-18, which coincides with a Saturday. This improvement is likely due to lower traffic demand on weekends, resulting in shorter queues that are easier to predict.
We visualized the estimation results for 2023-11-14 in Figure 8. The results indicate that both baselines are highly sensitive to errors in the aFCD data. As shown in Figure 8(a), these errors manifest in several ways: (a) during the congested period (15:00–19:00), relatively high speeds are recorded within the queue; (b) the first FCD segment (closest to the stop line) often reports higher speeds despite the presence of long queues, likely due to sparse probe vehicle coverage or smoothing artifacts introduced during aggregation; and (c) isolated low-speed patches appear in the middle of the segment, which do not align with the actual queue positions observed in the ground truth.
These inconsistencies in the aFCD signal lead to degraded performance in the baseline methods. Baseline OSD exhibits sharp, erratic fluctuations and fails to capture the afternoon peak queue. Baseline ISC, which relies on iso-speed contours, suffers from discontinuities and generates multiple queue estimates for a single time step, highlighting its sensitivity to noisy input data. In contrast, Q-Net successfully captures both the trend and variations in queue length, even under noisy input conditions, thanks to KalmanNet’s ability to dynamically track noise covariance. Furthermore, when we focus on the period from 13:30 to 14:00, we observe a queue formation delay in the aFCD. Despite this, Q-Net accurately tracks the onset of the queue, likely due to the initial estimate derived from the loop detector cumulative count data, which helps mitigate the lag present in the FCD signal.
6.3 Transferability
This subsection focuses on the transferability analysis of our method. We introduce two additional sections from the Rotterdam urban area that are adjacent to N1-IN: N3-IN and N3-OUT, as shown in Figure 9. Both sections are located underground and feature an above-ground split after vehicles exit their upstream intersections. These diverging flows are eventually merged at the downstream intersection. The upstream intersection of N3-IN includes lanes with combined directions (through and right-turn), meaning unobserved traffic flow can occur not only at the split but also through vehicles exiting in the studied direction.
Figure 9: Satellite view of N3-IN and N3-OUT. In both cases, the downstream area divides into an underground study section and an above-ground branch, which merge at the downstream intersection.
Previously, we evaluated our method on test data from the same section used for training in Subsection 6.2, demonstrating temporal transferability with spatial consistency. This scenario applies when sensors such as radars can be temporarily installed for training data collection and then removed to reduce long-term maintenance costs. Here, we investigate spatial transferability by training on one section (e.g., N1-IN) and testing on test days of different sections (e.g., N3-IN) without fine-tuning. This reflects scenarios where radar installation is infeasible or costly on adjacent sections but queue estimation remains necessary. Our adaptive design (Subsection 5.2) enables model transfer across sections of varying lengths.
Figure 10: Box plots of absolute estimation errors across all test days using different training sections.
(a) Model trained on N1-IN tested on N3-IN on 2023-11-14.
(b) Q-Net trained on N1-IN tested on N3-OUT..
Figure 11: Queue length estimation for section N3-IN and N3-OUT on 2023-11-14 using a model trained on N1-IN without fine-tuning.
Figure 10 presents absolute estimation errors under three training strategies: models trained individually on N1-IN, N3-IN, and N3-OUT, evaluated across all sections. As expected, each section achieves best performance when the model is trained on the same section, confirming effective capture of location-specific traffic dynamics. Performance on N3-IN and N3-OUT is similar regardless of which of these two sections is used for training, attributable to their spatial proximity and shared characteristics.
Interestingly, the model trained on N1-IN generalizes better to N3-IN and N3-OUT than vice versa. This asymmetry stems from N1-IN’s simpler traffic conditions and cleaner detector measurements, which enable learning of more accurate system dynamics representations with better convergence and stability. In contrast, N3-IN and N3-OUT exhibit higher complexity due to diverging over-ground branches that rejoin downstream (Figure 9). These branches create measurement challenges as downstream detector counts include vehicles not part of the initial section inflow, introducing unobserved traffic flows at both boundaries. The increased noise and structural ambiguities reduce the ability of data-driven models to learn reliable patterns, particularly when training data is limited.
Finally, Figure 11 visualizes the performance on N3-IN and N3-OUT for the test day 2023-11-14, using Q-Net trained on N1-IN. Our model successfully captures multiple queue peaks throughout the day, with the estimated queue (green line) closely following the ground truth (blue line) during most queue formation and dissipation phases. This demonstrates effective queue trend estimation for both unseen sections and time periods. Two noteworthy observations emerge from the results. First, the ground truth queue never drops to 0m, indicating potential radar malfunction since continuous queues even during early morning hours are unrealistic. Second, during afternoon peak periods when queues stabilize, Q-Net provides higher estimates than ground truth. This conservative bias likely reflects radar detection limitations and potential spillback to upstream intersections, but is acceptable since overestimation helps ensure activation of appropriate traffic control strategies.
6.4 Online estimation for real-time traffic control
Previously, in Subsection 4.1, we introduced the control input formulation based on queue length changes, which relies on estimating a fixed unobserved traffic flow rate via linear regression using cumulative net flow data over an entire day. While this approach is effective for ex post assessment and performance evaluation, it is not suitable for real-time applications where timely updates are essential. In this subsection, we extend our method to support real-time traffic control by dynamically updating the unobserved traffic flow rate during operation.
The underlying idea is straightforward: rather than computing a static unobserved traffic flow rate from full-day data, we continuously estimate it in real-time using cumulative flow data from the start of the day up to the current moment. Figure 12 compares the performance of the fixed (full-day) rate with the real-time adaptive version. Both methods are capable of capturing the general trend of queue dynamics. However, the real-time variant exhibits slightly reduced accuracy, which is expected given its reliance on shorter-term data that may be more susceptible to temporary fluctuations. Nonetheless, the ability to adapt online makes this approach more suitable for real-time traffic monitoring and control applications.
(a) Queue length estimation on 2023-11-14 (Tuesday).
(b) Queue length estimation on 2023-11-22 (Wednesday).
Figure 12: Comparison of queue length estimation for section N1-IN using two approaches for identifying unobserved traffic flow rate. The first estimates a fixed rate from full-day data via linear regression, while the second dynamically learns and adapts the rate in real time.
6.5 Ablation study
In this subsection, we examine the contributions of different components of our model through ablation analysis. Abewickrema et al. (2023) employed a Kalman filter for queue estimation, assuming that the maximum queue length remains constant over each signal cycle in the state evolution equation. This assumption motivates the design of our first variant, Q-Net w/o $u_{t}$, which excludes the control input derived from cumulative vehicle counts and instead assumes that the queue length remains unchanged over time. This variant relies solely on aFCD as input. We also introduce Q-EKF, in which the KalmanNet module is replaced by a standard Extended Kalman Filter (EKF), allowing us to isolate the benefit of using a learned observation update in the presence of nonlinear and partially known system dynamics.
Table 4: Performance comparison of model variants in the ablation study on section N1-IN.
All day Morning peak Afternoon peak
RMSE MAE MAPE RMSE MAE MAPE RMSE MAE MAPE
Default 71.65 49.51 48.60 5.29 36.94 79.43 26.79 68.04 17.91
Q-Net w/o $u_{t}$ 81.63 53.27 54.10 9.11 39.70 86.46 16.02 59.79 16.95
Q-EKF 229.79 189.27 292.12 175.09 180.39 446.28 117.27 121.00 49.83
Figure 13: Queue length estimation results for different model variants on 2023-11-14 for section N1-IN.
The performance comparison of the model variants is presented in Table 4. The default Q-Net demonstrates the best overall accuracy and outperforms all variants during the morning peak period. However, its performance is slightly worse than Q-Net w/o $u_{t}$ during the afternoon peak hours. This is likely because afternoon peaks are typically characterized by consistently long queues with minimal variation, reducing the benefit of incorporating queue change dynamics. Additionally, Q-EKF yields significantly higher errors, underscoring the importance of a learnable Kalman gain. This result highlights the advantage of using a data-driven approach to dynamically track error covariances, as implemented in KalmanNet.
In Figure 13, we visualize the queue length estimation results on 2023-11-14 produced by the model variants and our default configuration. Q-Net w/o $u_{t}$ exhibits a noticeable delay in capturing the queue formation during the morning peak period. This lag is likely due to noise in the aFCD data, which becomes more impactful in the absence of control input. Q-EKF, while able to follow some general trends, fails to accurately capture both the queue formation and dissipation phases. These findings underscore the limitations of fixed-gain approaches like EKF in the context of highly dynamic urban traffic, and collectively support the design choices underlying our proposed Q-Net.
7 Conclusion
In this study, we proposed Q-Net, a computationally and data-efficient framework for queue length estimation. Our problem setting reflects a general and practically relevant scenario in which traffic flow is only partially observed. Moreover, Q-Net relies solely on two widely available and privacy-preserving data sources: cumulative vehicle counts from loop detectors and aggregated floating car data (aFCD). Together, these aspects highlight the broad applicability of the proposed method.
The filtering algorithm of Q-Net leverages KalmanNet, an AI-augmented Kalman filter that combines the representational power of deep learning with the structure of state-space modeling. Unlike traditional Kalman filters, KalmanNet does not require prior knowledge of system or noise dynamics; instead, it learns to adapt dynamically to real-world noise patterns. Our contributions extend beyond a direct application of KalmanNet in three key ways: (a) Q-Net introduces a tailored state-space formulation specifically designed for queue length estimation. This formulation effectively fuses heterogeneous data sources with differing spatial and temporal resolutions. (b) Q-Net adapts KalmanNet to decouple measurement dimensionality from segment length, thereby enabling spatial transferability across different road segments. (c) Q-Net employs a training strategy designed to handle long time series data, ensuring stable and efficient model convergence.
Q-Net was implemented and evaluated on multiple road sections in Rotterdam, the Netherlands. Empirical results show that it consistently outperforms baseline methods during both the formation and dissipation phases of traffic queues, effectively correcting the lag typically observed in aFCD-based estimates. The model demonstrated strong temporal and spatial transferability, underscoring its practicality for real-world deployment without the need to maintain or install costly radar infrastructure. Furthermore, a real-time variant of Q-Net was introduced, illustrating its potential for dynamic traffic control applications.
Future research could enhance Q-Net by integrating graph machine learning to capture spatial correlations in road networks with multiple intersections, enabling more expressive evolution models for improved queue change estimation. Integrating additional contextual features, such as traffic signal timings or weather conditions, may further boost accuracy. Moreover, exploring alternative training paradigms, such as curriculum learning, and validating the approach across a more diverse set of urban environments would strengthen generalizability. Ultimately, combining Q-Net with traffic control and optimization strategies opens new avenues for intelligent, data-driven traffic management systems.